# Quantum Down Sampling Filter for Variational Auto-encoder

### v2

@@ 6.1 Reconstructed Images:
Figure 9 displays reconstructed grayscale digit images from the MNIST dataset, arranged in a grid format. The input image is first downscaled from 28x28 to 16x16, and then upscaled to 32x32. This sequence reveals that some images lose details during resolution reduction, leading to inaccuracies in their reconstruction. For example, the digit ’3’ in the 4th row and 1st column is successfully reconstructed by the Quantum VAE (Q-VAE), whereas the Classical VAE (C-VAE) and Direct Passing VAE (CDP-VAE) struggle to generate a well-defined version. However, in the 5th row and 3rd column, the digit ’0’ shows noticeable noise due to the resolution reduction, making it difficult to reconstruct accurately. In this case, C-VAE incorrectly generates a ’7’, while Q-VAE produces a hybrid image that blends ’7’ and ’0’. Most digits retain their recognizable shapes, but some exhibit blurriness or distortions as column 3, row 5 and ’6’ its difficult to recognise the image as ’7’ and ’9’, indicating challenges in capturing finer details during the reconstruction process. These issues may stem from factors such as limited latent space resolution’s used for the model (i.e 16), preprocessing steps like resizing or normalization, or suboptimal model architecture. Interestingly, prior evaluations using metrics like the FID have demonstrated that quantum-based approaches, particularly Quantum Variational Autoencoders (Q-VAE), achieve superior image quality on MNIST. The observed FID scores confirm that quantum-enhanced and hybrid methods produce sharper and more detailed reconstructions compared to classical models.
@@
Figure 9: Comparing Classical and Quantum Reconstructed Images. (a) MNIST Input First Down-sampled from 28x28 to 16x16 then Up-sampled Image from 16x16 to 32x32 resolution (b) Classical VAE Image Reconstructions, (c) MNIST CDP-VAE reconstructed images and (d) MNIST Q-VAE reconstructed images for Variational Autoencoder.
USPS images reconstructed from 3 different VAE models have been shown in Figure 10.
@@
Figure 10: Comparing Classical and Quantum Reconstructed Images. (a) USPS Input Up-sampled Image from 16x16 to 32x32 resolution (b) Classical VAE Image Reconstructions, (c) USPS CDP-VAE reconstructed images and (d) USPS Q-VAE reconstructed images for Variational Autoencoder.
@@
While this score indicates that the model can generate images with reasonable fidelity (i.e., the generated images resemble real USPS digits), it lags behind the performance of the Q-VAE. The quantum model’s ability to leverage more complex transformations in the latent space for image reconstruction enables it to capture more intricate patterns and feature representations in the data. This allows the Q-VAE to achieve lower FID scores (indicating higher image fidelity) compared to the CDP-VAE and C-VAE as seen in the results for the USPS dataset. For MSE, the Q-VAE again outperformed the other models, achieving a score of 0.0058. This lower error indicates that the quantum model is particularly effective at capturing the more complex, medium-resolution patterns in the USPS images. The Q-VAE’s quantum-enhanced encoding allows it to extract intricate features from the grayscale images, leading to more accurate reconstructions. Its performance on USPS highlights the potential of quantum circuits in improving image generation tasks, especially for datasets with more complex structures compared to MNIST. The C-VAE, with an MSE of 0.0070, performed worse than the Q-VAE and CDP-VAE on USPS, reflecting the same trend observed in MNIST. While still a competitive model, the C-VAE’s ability to reconstruct images is limited by its classical encoding mechanism. The model performs well, but its lack of advanced feature extraction techniques prevents it from achieving the same reconstruction quality as the Q-VAE. The CDP-VAE with an MSE of 0.0065, performance is better than the C-VAE on MNIST, it still lags behind both the Q-VAE on USPS. While reconstructed images we can see Q-VAE is reconstructing very well competing with C-VAE and CDP-VAE.
6.2 New Generated Images:
@@
The performance of the three models in handling random noise varies considerably. The generated images presented were captured at epoch 200. The Classical VAE requires more time to converge, necessitating longer training. Therefore, we evaluate the model at epoch 200 to assess its performance after sufficient training. While initial epochs Q-VAE has shown fast convergence at epoch 20, while other models were unable to generate digits even at epoch 20 .The Q-VAE outperforms the other models, achieving an FID score of 78.7 at epoch 200. This suggests that the Q-VAE is more resilient to noise, as it is capable of generating better-quality images even when starting from random noise inputs. In contrast, the C-VAE struggles with random noise, resulting in a much higher FID score of 94.4, indicating poor performance in reconstructing images from noisy inputs. The CDP-VAE also faces challenges with noise but performs slightly better than the C-VAE, with an FID score of 93.3. While both the C-VAE and CDP-VAE show limitations in noise handling, the Q-VAE demonstrates a notable advantage in generating more realistic images in the presence of noise. MNIST generated images from noise, have been shown below in Figure 11:
@@
Figure 11: MNIST New Generated Images from Random Noise
Figure 11 shows newly generated grayscale digit images, likely synthesized from random noise using VAE. These images are organized in a grid format, with varying levels of clarity and detail. While some digits are clearly defined and resemble their expected shapes, others appear blurry, distorted, or lack fine structural details, indicating limitations in the model’s ability to fully capture the underlying data distribution. These artifacts could result from constraints such as an insufficiently expressive latent space, challenges in training stability, or limitations in the model architecture. For instance, models with lower-resolution latent spaces may struggle to encode the nuanced variations required to generate sharp, realistic images.
@@
Figure 12: USPS New Generated Images from Random Noise
Interestingly, Q-VAE have demonstrated better performance in similar tasks, as reflected in improved FID scores. Q-VAE, in particular, have shown a remarkable ability to generate better-quality images from noise These operations are highly effective at capturing complex patterns and correlations in medium-resolution data, enabling more detailed and coherent outputs. Figure 12 shows how well Quantum can generate new images from noise while using only quantum superposition. The quantum advantage enables the model to effectively capture complex patterns from random noise, producing high-quality, leading to more realistic images [17] This highlights the advantage of quantum-based encoding in generating better-quality images, particularly in tasks that require nuanced and medium-resolution data representations, such as image reconstruction or generation.
@@ 6.2.1 Trainable Parameters
The results indicates that our proposed model Q-VAE and its similar approach in classical, CDP-VAE, both have shown significant improvement over C-VAE architectures in terms of parameter efficiency. Specifically, the C-VAE has a total of 407,377 trainable parameters, with 299,424 parameters dedicated to the encoder and 107,953 parameters assigned to the decoder. In contrast, the Quantum VAE significantly reduces the number of trainable parameters, totaling only 144,977 parameters. This is achieved by the quantum model’s more compact encoder, which contains only 37,024 parameters, a dramatic reduction compared to the classical VAE. This reduction highlights the efficiency of the quantum approach in learning and representing complex features with a smaller number of parameters, enabling better latent space compression.
Despite this reduction in parameters, the decoder structure remains similar between the quantum and classical models, both utilizing 107,953 parameters. This consistency in the decoder ensures that the Quantum VAE maintains comparable image reconstruction capabilities to the classical model, while benefiting from the quantum encoder’s enhanced efficiency. The lower parameter count in the QVAE encoder illustrates the potential of quantum circuits to offer a more resource-efficient alternative for training and generating better-quality image reconstructions. These findings underscore the practical benefits of quantum computing in optimizing VAE architectures, especially in tasks that require high-fidelity image generation and accurate reconstruction. While classical VAEs remain a strong choice for many machine learning applications, the Q-VAE shows promising advancements, particularly in fields where image quality and computational efficiency are paramount. Potential applications of quantum VAEs include digital imaging, healthcare diagnostics, and creative arts, where high-quality image generation is critical. However, despite these advantages, the adoption and effectiveness of Q-VAE still face challenges, including issues related to the scalability of quantum hardware and the complexity of quantum algorithms. Continued research and development in quantum machine learning are essential for unlocking the full potential of quantum-enhanced models, pushing the boundaries of what is achievable in generative modeling and real-world applications.
@@ 7 Conclusion
In this study, we introduce the Quantum Down-sampling Filter for Variational Auto-encoder (Q-VAE), a novel model that integrates quantum computing into the encoder component of a traditional VAE. Our experimental results demonstrate that Q-VAE offers significant advantages over classical VAE architectures across various performance metrics. In particular, Q-VAE consistently outperforms classical VAE, achieving lower FID scores, indicating superior image fidelity and the ability to generate better-quality reconstructions. The quantum-enhanced encoder captures more intricate features of the images, leading to improved generative performance. We evaluated the model on digit datasets such as MNIST and USPS, initially up-scaled to 32x32 for consistency. For a fair comparison, we also tested our quantum-enhanced model alongside a classical counterpart and a model mimicking classical behavior, referred to as Classical Direct Passing (CDP-VAE). One more key finding of this study is the significant reduction in the number of trainable parameters in the Q-VAE’s and CDP-VAE’s encoder. With only 37,024 parameters, the Q-VAE dramatically reduces computational complexity compared to the classical VAE, which contains over 407,000 trainable parameters. This reduction highlights the efficiency of quantum-based encoding methods, which offers the potential for more scalable and resource-efficient models. The results highlight the practical advantages of incorporating quantum computing into VAEs, particularly for tasks that require high-quality image generation and feature extraction. The Q-VAE not only improves image quality, but also enhances computational efficiency, making it a promising candidate for future applications in areas such as computer vision, data synthesis, and beyond. These findings suggest that further exploration of quantum-enhanced machine learning techniques could unlock new possibilities in generative modeling and image reconstruction, offering both performance improvements and computational savings.

### v3

@@ 6.2 Reconstructed Images
Figure 12 displays reconstructed grayscale digit images from the MNIST dataset, arranged in a grid format. The input image is first downscaled from 28x28 to 16x16, and then upscaled to 32x32. This sequence reveals that some images lose details during resolution reduction, leading to inaccuracies in their reconstruction. For example, the digit ’3’ in the 4th row and 1st column is successfully reconstructed by Q-VAE, whereas the C-VAE and CDP-VAE struggle to generate a well-defined version. However, in the 5th row and 3rd column, the digit ’0’ shows noticeable noise due to the resolution reduction, making it difficult to reconstruct accurately. In this case, C-VAE incorrectly generates a ’7’, while Q-VAE produces a hybrid image that blends ’7’ and ’0’. Most digits retain their recognizable shapes, but some exhibit blurriness or distortions as column 3, row 5 and ’6’ its difficult to recognise the image as ’7’ and ’9’, indicating challenges in capturing finer details during the reconstruction process. These issues may stem from factors such as limited latent space resolution’s used for the model (i.e 16), preprocessing steps like resizing or normalization, or suboptimal model architecture. Interestingly, prior evaluations using metrics like the FID have demonstrated our proposed model Q-VAE, achieve better image quality on MNIST.
@@
Figure 12: Comparing classical and quantum reconstructed images. (a) MNIST input image (b) C-VAE image reconstructions, (c) CDP-VAE image reconstruction and (d) Q-VAE image reconstruction.
On the MNIST dataset, the Q-VAE achieved the lowest MSE of 0.0088, demonstrating its ability to produce high-fidelity reconstructions compared to the other models. The quantum-based encoding method enables the Q-VAE to extract richer feature representations, leading to more accurate reconstructions. The quantum model’s ability to capture complex dependencies between pixels allows it to outperform classical approaches, even when dealing with highly structured images like those in the MNIST dataset. The C-VAE, with an MSE of 0.0093, showed slightly higher error compared to the Q-VAE, indicating that while it performs well, its feature extraction capabilities are not as advanced as those of the quantum model. Despite its relatively low MSE, the classical model’s performance is constrained by the limitations of traditional neural network-based encoding and decoding methods, which cannot capture the full complexity of the data as efficiently as quantum circuits. The CDP-VAE, with an MSE of 0.0092, performs little better then C-VAE but still didn’t perform better then Q-VAE as shown in Figure 12.
@@
Figure 13: Comparing classical and quantum reconstructed images. (a) USPS input up-sampled image from 16x16 to 32x32 resolution (b) USPS C-VAE image reconstructions, (c) USPS CDP-VAE reconstructed images and (d) USPS Q-VAE reconstructed images.
@@
USPS images reconstructed from 3 different VAE models have been shown in Figure 13. For the USPS dataset, while FID score for C-VAE indicate that the model can generate images with reasonable fidelity (i.e., the generated images resemble real USPS digits), it lags behind the performance of the Q-VAE. The quantum model’s ability to leverage more complex transformations in the latent space for image reconstruction enables it to capture more intricate patterns and feature representations in the data. This allows the Q-VAE to achieve lower FID scores (indicating higher image fidelity) compared to the CDP-VAE and C-VAE as seen in the results for the USPS dataset. For MSE, the Q-VAE again outperformed the other models, achieving a score of 0.0058. This lower error indicates that the quantum model is particularly effective at capturing the more complex, medium-resolution patterns in the USPS images. The Q-VAE’s quantum-enhanced encoding allows it to extract intricate features from the grayscale images, leading to more accurate reconstructions. Its performance on USPS highlights the potential of quantum circuits in improving image generation tasks, especially for datasets with more complex structures compared to MNIST. The C-VAE, with an MSE of 0.0070, performed worse than the CDP-VAE and Q-VAE on USPS, reflecting the same trend observed in MNIST. While still a competitive model, the C-VAE’s ability to reconstruct images is limited by its classical encoding mechanism. The model performs well, but its lack of advanced feature extraction techniques prevents it from achieving the same reconstruction quality as the Q-VAE. The CDP-VAE with an MSE of 0.0065, performance is better than the C-VAE on MNIST which is 0.0092, it still lags behind both the CDP-VAE and Q-VAE on USPS. While reconstructed images we can see Q-VAE is reconstructing very well competing with C-VAE and CDP-VAE.
6.3 New Generated Images
@@
The performance of the three models in handling random noise varies considerably. The generated images presented were captured at epoch 200. The C-VAE requires more time to converge, necessitating longer training. Therefore, we evaluate the model at epoch 200 to assess its performance after sufficient training. While initial epochs Q-VAE has shown fast convergence at epoch 20, while other models were unable to generate digits even at epoch 20. The Q-VAE outperforms the other models, achieving an FID score of 78.7 at epoch 200. In contrast, the C-VAE struggles with random noise, resulting in a much higher FID score of 94.4, indicating poor performance in generating from latent space using Gaussian variable as input. The CDP-VAE also faces challenges with noise but performs slightly better than the C-VAE, with an FID score of 93.3. While both the C-VAE and CDP-VAE show limitations in noise handling, the Q-VAE demonstrates a notable advantage in generating more realistic images from random variables. MNIST generated images from noise, have been shown below in Figure 14:
@@
Figure 14: MNIST image generation from Gaussian noise.
These images are organized in a grid format, with varying levels of clarity and detail. While some digits are clearly defined and resemble their expected shapes, others appear blurry, distorted, or lack fine structural details, indicating limitations in the model’s ability to fully capture the underlying data distribution. These artifacts could result from constraints such as an insufficiently expressive latent space, challenges in training stability, or limitations in the model architecture. For instance, models with lower-resolution latent spaces may struggle to encode the nuanced variations required to generate sharp, realistic images.
@@
Figure 15: USPS image generation from Gaussian noise.
Interestingly, Q-VAE have demonstrated better performance in generating images from noise, as reflected in improved FID scores. These operations are highly effective at capturing complex patterns and correlations in medium-resolution data, enabling more detailed and coherent outputs. Similar results were obtained for USPS as shown in Figure 15.
@@ 6.3.1 Trainable Parameters
The results indicate that both the proposed Q-VAE model and the CDP-VAE demonstrate significant improvements over the C-VAE architecture in terms of parameter efficiency. Specifically, the C-VAE consists of 407,377 trainable parameters, with 299,424 allocated to the encoder and 107,953 to the decoder. In contrast, the CDP-VAE reduces the total number of trainable parameters to 144,977, primarily due to its more compact encoder, which contains only 37,024 parameters. While the Q-VAE also has 144,977 parameters, this reduction is attributed to the CDP-VAE’s design rather than the quantum encoding itself. This highlights the role of CDP-VAE in achieving parameter efficiency while leveraging quantum encoding for enhanced feature extraction. Despite the reduction in parameters, the decoder structure remains unchanged across all models, with 107,953 parameters. Notably, the reduction in trainable parameters stems from CDP-VAE, not the quantum component, as the quantum circuit in Q-VAE does not introduce additional trainable parameters. As a result, both CDP-VAE and Q-VAE share the same total number of parameters, totaling 144,977. This demonstrates the efficiency of the CDP approach in reducing parameter overhead while maintaining effective feature extraction.
These findings highlight the role of CDP-VAE in optimizing parameter efficiency within VAE architectures while ensuring high-fidelity image reconstruction. Although C-VAEs remain a widely adopted choice in machine learning, the combination of CDP-VAE’s parameter reduction and Q-VAE’s quantum encoding presents a compelling direction for future research. Potential applications include digital imaging, healthcare diagnostics, and creative arts, where both high-quality image generation and computational efficiency are crucial. However, the effectiveness and adoption of Q-VAE remain subject to the challenges of quantum hardware scalability and the complexity of quantum algorithms. Continued advancements in quantum machine learning are essential to further explore the integration of quantum encoding in generative models, paving the way for more efficient and powerful solutions in real-world applications.
However, despite these advantages, the adoption and effectiveness of Q-VAE still face challenges, including issues related to the scalability of quantum hardware and the complexity of quantum algorithms. Continued research and development in quantum machine learning are essential for unlocking the full potential of quantum-enhanced models, pushing the boundaries of what is achievable in generative modeling and real-world applications.
@@ 7 Conclusion
In this study, we introduce the quantum down sampling filter for Q-VAE, a novel model that integrates quantum computing into the encoder component of a traditional VAE. Our experimental results demonstrate that Q-VAE offers significant advantages over classical VAE architectures across various performance metrics. In particular, Q-VAE consistently outperforms C-VAE, achieving lower FID scores, indicating superior image fidelity and the ability to generate better-quality reconstructions. The quantum-enhanced encoder captures more intricate features of the images, leading to improved generative performance. We evaluated the model on digit datasets such as MNIST and USPS, initially up-scaled to 32x32 for consistency. For a fair comparison, we also tested our quantum-enhanced model alongside a classical counterpart and a model mimicking classical behavior, referred to as CDP-VAE. One key finding of this study is the significant reduction in the number of trainable parameters in the CDP-VAE’s encoder. While the C-VAE consists of over 407,000 trainable parameters, the CDP-VAE achieves a reduction by significantly lowering the number of parameters. Furthermore, the Q-VAE leverages quantum encoding, which does not add any additional trainable parameters. This combination highlights the efficiency of quantum-based encoding methods and the potential for more scalable and resource-efficient models without increasing computational complexity. The results highlight the practical advantages of incorporating quantum computing into VAEs, particularly for tasks that require high-quality image generation and feature extraction. The Q-VAE not only improves image quality, but also enhances computational efficiency, making it a promising candidate for future applications in areas such as computer vision, data synthesis, and beyond. These findings suggest that further exploration of quantum-enhanced machine learning techniques could unlock new possibilities in generative modeling and image reconstruction, offering both performance improvements and computational savings.
Data Availability: The data is available at https://github.com/RRFar/Quantum-Downsampling-Filter.
Conflict of Interest: The authors declare no conflict of interest.